Following is the information of two plants located at ABC City and XYZ City

ABC City Plant
Total Fixed Cost (TFC) $100,000
Average Cost per unit (AVC) $10
Selling Price per unit $20
Capital labor (K/L) 2:1
Output rate (Q) 22,000

XYZ City Plant

Total Fixed Cost (TFC) $200,000
Average Cost per unit (AVC) $8
Selling Price per unit $22
Capital labor (K/L) 2:1
Output rate (Q) 16,000

Determine the profit elasticity for each plant.
1
Expert's answer
2015-12-28T12:51:37-0500
profit=TR-TC=P*Q-AVC*Q-TFC=Q(P-AVC)-TFC
For ABC Plant profit is equal (20-10)*22000-100000=120000
For XYZ Plant it is equal (22-8)*16000-200000=24000

The elasticity of profit is equal (MR-MC) *Q/profit.
For ABC Plant it is equal (20-10)*22000/120000=11/6
For XYZ Plant it is equal (22-8)*16000/24000=28/3